whats the value of algebra problem if g(x)=x2-2x g(1/5)
Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(32388)
Answer by ewatrrr(24785)


Hi
If I am understanding the queston properly
g(x)=x^2-2x
g(1/5) Substituting x =(1/5)
g(1/5) = 1/25 - 2/5 = 1/25 - 10/25 = (-9/25)
g(1/5)= (-9/25)